# Leverage Knowledge Graph and Large Language Model for Law Article Recommendation: A Case Study of Chinese Criminal Law

### v1

@@ 3.4.1 Analyzing the Impact of Label Imbalance in Training
We utilized the BERT model for law article recommendation as comparison methods. This method was compared against the method proposed in this paper: the LLM(specifically OpenAI’s ChatGPT-4.0) for law article recommendation based on CLAKG (with case information).
Table 5: Model Accuracy Comparison Between Bert and Ours
@@
LLM+CLAKG (OURS, with case) 0.694
As shown in the Table 5, the accuracy of the BERT model is only 0.289. Upon analyzing the results provided by the BERT method, we discovered that this method classified all test cases under ”Article 133.” We believe this could be due to the small size of the training set, an imbalance in training labels [42], or an inappropriate loss function. Even after replacing the cross-entropy loss function with the Focal Loss function, the model still classified all test cases as ”Article 133”.
The proposed model (LLM+CLAKG) achieved an accuracy of 0.694, which is significantly higher than that of the BERT-based method. This indicates that the proposed approach effectively mitigates the impact of the insufficient data and label imbalance in training.
@@ 3.4.2 Analyzing the Impact of Hallucinations in LLMs
We utilized the LLM (specifically OpenAI’s ChatGPT-4.0, as well as other LLMs used in different methods) for law article recommendation as comparison method. This method was compared against the method proposed in this paper: the LLM for law article recommendation based on CLAKG (with case information).
Table 6: Model Accuracy Comparison Between LLM and Ours
@@
As shown in the Table 6, the accuracy of the LLM without external legal data is significantly higher than that of the BERT model, achieving an accuracy of 0.549. However, it is still lower than the proposed model (LLM+CLAKG). We attribute this to the hallucinations of LLM and the proposed approach effectively mitigates the impact of label imbalance in training. The proposed model (LLM+CLAKG) achieved an accuracy of 0.694, which is significantly higher than that of the LLM-based method. This indicates that the proposed approach effectively mitigates the Huallucinations in LLMs.
3.4.3 Analyzing the Impact of Utilizing Macro Semantics
We utilized the LLM (specifically OpenAI’s ChatGPT-4.0, as well as other LLMs used in different methods) based on TFIDF for law article recommendation as comparison method. This method was compared against two method proposed in this paper: the LLM for law article recommendation based on CLAKG (with case information) and the LLM for law article recommendation based on CLAKG (no case information).
Table 7: Model Accuracy Comparison Between LLM+TFIDF and Ours
Model Accuracy
LLM+TFIDF (base line) 0.595
LLM+CLAKG (OURS, no case) 0.676
LLM+CLAKG (OURS, with case) 0.694
As shown in the Table 7, both the TFIDF-based law article recommendation method and the CLAKG-based law article recommendation method mitigate the hallucinations of the LLM and improve its accuracy, with the latter performing better. We believe this is because the former only matches at the text level, while the latter matches the new case with keywords at the semantic level. Additionally, the latter utilizes all information within the CLAKG through graph embedding techniques, allowing it to calculate the most relevant law articles from a holistic perspective.

### v2

@@ 4.4.1 Analyzing the Impact of Label Imbalance in Training
We utilized the BERT[23], GRU[8] and DPCNN[11] model for law article recommendation as comparison methods. These methods were compared against the method proposed in this paper: the LLM(specifically OpenAI’s ChatGPT-4.0) for law article recommendation based on CLAKG (with case information).
Table 5: Model accuracy comparison between classical models and ours
@@
GRU (baseline) 0.427
DPCNN (baseline) 0.495
LLM + CLAKG (OURS, no case) 0.676
As shown in the Table 5, the accuracy of the BERT (baseline) is only 0.289. Upon analyzing the results provided by the BERT method, we discovered that this method classified all test cases under ”Article 133.” We believe this could be due to the small size of the training set, an imbalance in training labels[40], or an inappropriate loss function. Even after replacing the cross-entropy loss function with the Focal Loss function, the model still classified all test cases as ”Article 133”.
The GRU (baseline) model achieved an accuracy of 0.427. GRU, or Gated Recurrent Unit, is a type of recurrent neural network (RNN) that is widely used for sequential data, and this result represents its performance on the task at hand.
The DPCNN (baseline) model, which stands for Deep Pyramid Convolutional Neural Network, performed slightly better, with an accuracy of 0.495. DPCNN is a convolutional neural network architecture designed for text classification tasks, and its accuracy reflects its ability to capture hierarchical patterns in the input data.
The proposed model (LLM+CLAKG) achieved an accuracy of 0.694, which is significantly higher than any one of the classic methods. This indicates that the proposed approach effectively mitigates the impact of the insufficient data and label imbalance in training.
@@ 4.4.2 Analyzing the Impact of Hallucinations in LLMs
We utilized the LLM[41] (specifically OpenAI’s ChatGPT-4.0, as well as other LLMs used in different methods), TFIDF-RAG[21], Graph-RAG[27] and Light-RAG[28] for law article recommendation as comparison method. This method was compared against the method proposed in this paper: the LLM for law article recommendation based on CLAKG (without case information).
Table 6: Comparing proposed method with other RAG method
@@
TFIDF-RAG (baseline) 0.597
Graph-RAG (baseline) 0.179
Light-RAG(baseline) 0.636
LLM+CLAKG (OURS, no case) 0.676
As shown in the Table 6, the accuracy of the LLM without external legal data is significantly higher than that of the classic model, achieving an accuracy of 0.549. Graph-RAG’s extremely low result may be due to its graph construction or retrieval strategy not aligning well with our dataset, indicating challenges in directly applying that method to this legal domain. The proposed model (LLM+CLAKG) achieved an accuracy of 0.676, which is significantly higher than that of the TFIDF-RAG, Light-RAG, Graph-RAG method. This indicates that the proposed approach effectively mitigates the hallucinations in LLMs. We believe this is because other RAG method only matches at the text level, while the latter matches the new case with keywords at the semantic level. Additionally, the latter utilizes all information within the CLAKG through graph embedding techniques, allowing it to calculate the most relevant law articles from a holistic perspective.
4.4.3 Ablation Study
In this ablation study, we evaluate the impact of incorporating different types of information on the performance of LLMs for law article recommendation tasks. Specifically, we compare the performance of the baseline models with and without additional case and law article information.
We conducted experiments using both the proprietary GPT-4 model and the open-source Llama-3.1-405b-chat model. The baseline results for GPT-4 showed an accuracy of 0.549, while incorporating case-specific information with CLAKG (Case Law and Statutory Knowledge Graph) improved the accuracy to 0.694. Similarly, for the Llama-3.1-405b-chat model, the baseline performance was 0.619, and adding reference case and statutory information boosted the accuracy to 0.786.
These results clearly demonstrate that the inclusion of relevant information, such as reference cases, significantly enhances model performance across both proprietary and open-source models. This validates the robustness and versatility of our approach, showing its applicability not only to high-performing proprietary models but also to widely available open-source models. The detailed results are summarized in Table 7.
Table 7: Results of different models in the ablation study.
Model Accuracy
LLM (GPT-4, baseline) 0.549
@@ 4.4.3 Ablation Study
LLM (llama-3.1-405b-chat, baseline) 0.619
LLM + CLAKG (llama-3.1-405b-chat, with case) 0.786
This ablation study highlights the significant impact that additional legal context, such as case and law article information, can have on the performance of LLMs for law article recommendation tasks, confirming the effectiveness of our proposed methodology.